# CHIPS-FF: Evaluating Universal Machine Learning Force Fields for Material Properties

### v3

@@ Abstract
A majority of the reference data used to compare our uMLFF results to were obtained from the JARVIS-DFT database, which contains relaxed structural data (primitive lattice vectors, volume), elastic properties (bulk modulus, elastic tensor) and phonon band structure 38, 39. Within JARVIS-DFT, there exists a database of single vacancy calculations 16 and surfaces 21 that were both used for benchmarking these uMLFFs. All of the DFT data in JARVIS was computed with the vdW-DF-optB88 111 functional, which is a modified Generalized Gradient Aprroximation (GGA) functional designed to explicitly include van der Waals (vdW) interactions. This is slightly different than other materials repositories such as the Materials Project (MPtrj), which mostly contains DFT calculations performed with the Perdew Burke Ernzerhof (PBE) 112 GGA functional. Despite mostly comparing to JARVIS-DFT results (vdW-DF-optB88), the CHIPS-FF workflow has a flexible framework to use any DFT dataset as a ground truth (i.e., the Materials Project, Alexandria, a users own DFT dataset, etc.). In addition, the platform is flexible enough to compare uMLFF results directly to available experimental values.
@@
uMLFF Type Bulk Surface Vacancy alignn_ff 6.7 37.5 39.6 chgnet 0.0 1.1 6.2 eqV2_153M_omat 0.0 14.8 31.2 eqV2_31M_omat 0.0 17.0 25.0 eqV2_31M_omat_mp_salex 0.0 5.7 0.0 eqV2_86M_omat 1.9 18.2 27.1 eqV2_86M_omat_mp_salex 0.0 5.7 2.1 mace 0.0 6.8 25.0 mace-alexandria 2.9 15.9 33.3 mace-mpa 0.0 6.8 33.3 matgl 0.0 4.5 14.6 matgl-direct 3.8 6.8 12.5 mattersim 0.0 4.5 16.7 orb-d3-v2 0.0 1.1 2.1 orb-v2 1.0 0.0 0.0 sevennet 0.0 2.3 41.7
@@
uMLFF Type err_a err_b err_c err_vol err_kv err_c11 err_c44 err_phonon Time Å Å Å Å3 GPa GPa GPa cm-1 seconds alignn_ff 0.087 0.102 0.137 9.95 122 196 74 150 8.8 x 104 chgnet 0.046 0.049 0.109 3.58 85 59 45 68 3.5 x 104 eqV2_153M_omat 0.036 0.040 0.111 2.90 117 70 32 49 1.6 x 102 eqV2_31M_omat 0.029 0.032 0.099 3.05 115 78 37 48 3.4 x 105 eqV2_31M_omat_mp_salex 0.028 0.030 0.096 3.17 112 47 41 48 2.7 x 105 eqV2_86M_omat 0.033 0.040 0.094 3.41 116 85 35 49 9.7 x 105 eqV2_86M_omat_mp_salex 0.027 0.030 0.100 3.14 114 45 35 48 7.2 x 105 mace 0.035 0.038 0.084 3.00 94 43 38 58 7.2 x 104 mace-alexandria 0.081 0.086 0.206 6.05 76 96 45 81 2.5 x 105 mace-mpa 0.044 0.046 0.098 3.57 110 35 34 49 8.4 x 104 matgl 0.052 0.057 0.128 3.50 83 64 41 78 1.7 x 104 matgl-direct 0.044 0.046 0.106 2.88 84 73 50 85 2.5 x 104 mattersim 0.031 0.033 0.110 3.18 110 32 32 48 5.9 x 104 orb-d3-v2 0.031 0.030 0.048 1.90 130 74 74 52 2.7 x 104 orb-v2 0.023 0.025 0.159 3.38 119 80 74 49 2.4 x 104 sevennet 0.035 0.038 0.100 3.43 103 46 37 53 5.7 x 104
@@
When considering the use of a pretrained uMLFF, the most important considerations should be accuracy, rate of convergence, scalability, and computational cost. After running the full CHIPS-FF workflow for the test set of 104 materials, we performed an analysis of which structural relaxation calculations reached convergence (force-maximum threshold of 0.05 eV/Å within 200 steps) for bulk materials, surfaces and vacancies. This analysis is presented in Table 1, where the percentage of unconverged results is depicted for each calculation type. Aside from ALIGNN-FF, all other uMLFF have a near-perfect convergence rate for bulk structures. With regards to surface and defect calculations, the convergence rate is more variable. Interestingly, the OMat24 models significantly vary in terms of convergence for surfaces and defects depending on training data. The OMat24 models trained on OMat+MPTrj+sAlexandria have a significantly higher rate of convergence when compared to the models solely trained on OMat. ORB and CHGNet have a successful rate of convergence compared to other models, which is remarkable due to the fact that the computational cost is so low (see Table 1). This is especially important since the computational cost of optimizing surfaces and defect structures greatly outweighs bulk unit cells.
For each of the 104 structures, we performed structural relaxations, fitting of the energy vs. volume curve (equation of state), calculation of the elastic tensor, calculation of the phonon band structure, relaxation of all nonpolar surfaces (among [1, 0, 0], [1, 1, 1], [1, 1, 0], [0, 1, 1], [0, 0, 1], and [0, 1, 0]) , and relaxation of each type of point defect in the material. The computational timing and errors in the lattice constants, elastic properties and phonons for each uMLFF model are depicted in Table 2. Out of the 104 materials in our test set, there were 85 matching entries for DFT phonon calculations in JARVIS, from which the composite MAE of the phonon band structure was computed in Table 2. For most structural and phonon properties, the OMat24 models perform exceedingly well, but are the most computationally intensive. Fitting the equation of state to obtain bulk modulus and calculating the elastic tensor prove to be a more challenging task for these uMLFF models, emphasizing difficulties in modeling the potential energy landscape far from equilibrium. We found that models such as MACE-Alexandria and MatGL have superior performance in predicting the energy vs. volume curves compared to the OMat and ORB models. For contributions to the elastic tensor (C11 and C44), we also observe relatively high errors. We see that only a few models can simultaneously predict C11 and C44 with reasonable accuracy (MAE less than 50 GPa). Interestingly, we find that eqV2_31M_omat, eqV2_86M_omat and eqV2_153M_omat can predict C44 with reasonable accuracy, but have much higher prediction errors for C11, while eqV2_31M_omat_mp_salex and eqV2_86M_omat_mp_salex do a better job at predicting C11 and C44 simultaneously. We observe that MACE-MPA-0 and MatterSim give the best results for simultaneous C11 and C44 predictions.
Excellent results for structural parameters and phonons are reported for orb-v2 and MatterSim at 10 - 100 times less the computational cost of the OMat24 models. ORB and MatterSim offer similar performance in terms of predicting phonon band structure when compared to several equivariant uMLFFs such as MACE, OMat24 and SevenNet. This is especially interesting since ORB and MatterSim-v1 are not equivariant uMLFFs. In addition, we see improved phonon results for the newer MACE-MPA-0 when compared to MACE-MP-0. Recently, it was reported that OMat and ORB models had substantial errors in phonon predictions when compared to models such as MatterSim, MACE, CHGNet, and SevenNet for 10,000 materials 90. It is entirely possible that ORB and OMat models can predict accurate phonons for our limited test set (104 materials) but fail for a broader and more diverse set of materials.
@@
One interesting trend we observe in our results for this set of 104 materials is that the error for lattice constant c is substantially higher than a and b. This can be due in part to the test set containing a significant portion ( 10 ) of vdW bonded (“2D-bulk-like”) layered structures (see Fig. S1). Since most of these uMLFF were trained on PBE data without vdW corrections (with the exception of ALIGNN-FF), we can expect larger errors for vdW systems. For example, we observe significant errors for hexagonal BN (JVASP-62940) when relaxed with orb-v2. This calculation did not converge within 200 steps and in the final step we observed over 130 error in c. Aside from this problematic structure (considering only the 103 structures that did converge during structural relaxation with ORB), the ORB MAE for c was 0.08 Å. Fortunately, some of these models such as MACE and ORB have explicit dispersion corrections that can be added to more accurately address vdW interactions 68, 69, 54, 55. To test this, we ran the workflow for dispersion corrected orb-d3-v2. For orb-d3-v2, all 104 bulk structures converged and we observe a significant reduction in error for c (0.048 Å) and voulme (1.90 Å3). For JVASP-62940 relaxed with orb-d3-v2, the percent error in c is 0.5 . This emphasizes the importance of dispersion corrections in uMLFF architectures for vdW bonded materials. An interactive version of these results is available on the JARVIS-Leaderboard platform (see Fig. 2b) for Pearson Correlation Coefficient metric), where it will be continuously updated as new uMLFF models are developed and released.
@@
In addition to summing the computational timing for the entire workflow, we also examined the scaling behavior of each uMLFF (on CPU). These results for supercells of Cu (up to 10,000 atoms) are depicted in Fig. 2a). Unsurprisingly, we see that OMat model scales much more drastically with system size as compared to the other uMLFF models. In addition, we find that the invariant models have more favorable scaling than the equivariant models such as MACE and SevenNet, with SevenNet having slightly better scaling than MACE. The scalability of ORB and MatterSim is a huge advantage due to their accuracy of various material property predictions. Similar scalability studies were conducted in Ref. 55 The quantities of surface energy and vacancy formation energy are difficult to predict with machine learning methods and provide a rigorous benchmarking test for uMLFF architectures. Out of the 104 materials, there exists 85 entries for surface DFT calculations and 48 entries for defect DFT calculations within JARVIS-DFT. We used these calculations to compare our uMLFF results. Fig. 3a) and Fig. 3b) depict the parity plots and corresponding MAE for surface energy and vacancy formation energy of our test sets. The very low error for surface energy (0.18 J/m2) and vacancy formation energy (0.94 eV) highlight some of the successes of recent state-of-the-art models such as OMat24, ORB, MACE-MPA-0 and MatterSim. We also observe generally good results for MACE-MP-0 and SevenNet when predicting vacancy formation energy and surface energies. The low cost of ORB, high rate of convergence and high accuracy make it a viable tool to relax larger surfaces and defect supercells.
@@
In addition to benchmarking the properties of bulk materials, we utilized the CHIPS-FF workflow for interface calculations. Fig. 3c depicts the parity plot for work of adhesion (Wad) for various material interfaces. The MAE for Wad is computed with respect to “ground truth” data taken from Ref. 113 (from experiment and theory). As seen in Fig. 3c, predictions for Wad are extremely poor for each uMLFF type. This is not surprising due to the fact that none of these uMLFF are trained on interface data. We went on to analyze one of the best performing predictions for an Al(111)Al2O3(001) interface (see red box on Fig. 3c). From here, we performed an in-plane (xy) grid search to find the optimal energy configuration of the interface. In addition, this in-plane scan can be a test of how smooth the potential energy surface is. On the right panel of Fig. 3c we see the results of this interface scan for a number of uMLFFs. We observe a smoother potential energy surface for MACE, SevenNet, MatterSim and OMat24 models. Although the MAE is extremely high for all uMLFF models for Wad, it is the lowest for ALIGNN-FF. This can be due in part to ALIGNN-FF being trained on vdW-corrected DFT calculations and having a substantial amount of exfoliable materials in the dataset.
@@
In addition to testing these uMLFF models on a smaller and more focused dataset, we also tested on a larger and more diverse dataset for force predictions. Unlike total energy, forces are more independent of exchange-correlation functional and the accuracy of force predictions on various datasets can give valuable insight into how a particular model performs. We chose to benchmark each uMLFF on the MLEARN dataset from Ref. 114, which consists of face-centered cubic (Cu, Ni) and body-centered cubic (Li, Mo) metals and diamond group IV semiconductors (Si, Ge) which span a wide range of crystal structures and bonding environments ( 200 - 300 data points for each element). These results are depicted in Table 3. In addition, we benchmarked the accuracy of force predictions on very large datasets that were used to train uMLFFs. These datasets included ALIGNN_FF_DB (307,000 used to train ALIGNN-FF), MPF (188,000 used to train M3GNet), and MPTrj (1.58 million used to train CHGNet, MACE, SevenNet and used in the training of ORB and OMat models). These results are depicted in Table 4. Unsurprisingly, superior accuracy is obtained across all datasets when the OMat and ORB models are used to predict forces. We also observe highly accurate forces for MatterSim and a substantial reduction in force error when going from MACE-MP-0 to MACE-MPA-0. The CHIPS-FF package has built in functions to perform these benchmarking calculations on each respective dataset, with a flexible framework to add additional datasets (such as OMat and Alexandria) and new uMLFFs as they develop over time.
@@
ALIGNN_FF_DB (307k) MPF (188k) MPTrj (1.58M) MLFF Type Error (eV/Å) Time (s) Error (eV/Å) Time (s) Error (eV/Å) Time (s) alignn_ff 0.472 2.9 x 104 0.639 4.1 x 104 0.581 3.5 x 105 chgnet 0.088 3.2 x 104 0.101 1.1 x 105 0.060 7.6 x 105 eqV2_153M_omat 0.050 1.5 x 106 - - - - eqV2_31M_omat 0.051 3.0 x 105 0.046 3.5 x 105 - - eqV2_31M_omat_mp_salex 0.055 3.9 x 105 0.032 4.3 x 105 - - eqV2_86M_omat 0.052 9.2 x 105 0.043 9.3 x 105 - - eqV2_86M_omat_mp_salex 0.056 9.2 x 105 - - - - mace 0.104 5.0 x 104 0.125 1.8 x 105 0.059 1.1 x 106 mace-alexandria 0.221 1.3 x 105 0.395 2.7 x 105 - - mace-mpa 0.072 7.0 x 105 0.061 1.3 x 105 - - matgl 0.091 1.1 x 104 0.066 3.4 x 104 0.100 1.4 x 105 matgl-direct 0.104 1.7 x 104 0.095 4.2 x 104 0.118 2.7 x 105 mattersim 0.071 8.7 x 104 0.047 7.1 x 104 0.059 1.7 x 106 orb-v2 0.062 5.0 x 104 0.049 8.4 x 104 0.028 7.5 x 105 sevennet 0.092 4.4 x 104 0.081 6.0 x 104 0.041 5.0 x 105

### v4

@@ Abstract
A majority of the reference data used to compare our uMLFF results to were obtained from the JARVIS-DFT database, which contains relaxed structural data (primitive lattice vectors, volume), elastic properties (bulk modulus, elastic tensor) and phonon band structure 37, 38. Within JARVIS-DFT, there exists a database of single vacancy calculations 15 and surfaces 20 that were both used for benchmarking these uMLFFs. All of the DFT data in JARVIS was computed with the vdW-DF-optB88 115 functional, which is a modified Generalized Gradient Aprroximation (GGA) functional designed to explicitly include van der Waals (vdW) interactions. This is slightly different than other materials repositories such as the Materials Project (MPtrj), which mostly contains DFT calculations performed with the Perdew Burke Ernzerhof (PBE) 116 GGA functional. Despite mostly comparing to JARVIS-DFT results (vdW-DF-optB88), the CHIPS-FF workflow has a flexible framework to use any DFT dataset as a ground truth (i.e., the Materials Project, Alexandria, a users own DFT dataset, etc.). In addition, the platform is flexible enough to compare uMLFF results directly to available experimental values. Of course, comparing the results of a quantity computed with a uMLFF to a DFT result with an arbitrary exchange-correlation functional and varying convergence criteria may result in biased or inconclusive error metrics. In this case, it may be suitable to compare uMLFF results directly to available experimental data, especially if the uMLFF prediction is closer to “reality” than the DFT calculation it is being evaluated against. Although it is possible to assess the accuracy of MLFFs by comparing results directly to DFT or experiment, there is a lack of explicit uncertainty quantification (UQ) for most MLFFs, including the universal pretrained MLFFs mentioned in this work. A robust UQ method should account for uncertainty arising from measurement noise (aleatoric) and uncertainty in predictions arising from model error (epistemic). Due to the fact that the DFT training data does not possess aleatoric error, most of the uncertainty for MLFFs is epistemic, which can be due to scarcity of data, limitations of model architecture, or poor parameter optimization within the model. Recently, different epistemic UQ methods such as ensemble-based uncertainty 117, deep evidential regression 118, mean-variance estimation 119, Gaussian mixture models 120 have been implemented and benchmarked for atomistic MLFFs 121, 122. It is evident that low-cost and widespread implementations of UQ for universal pretrained uMLFFs is a necessary future development.
@@
uMLFF Type Bulk Surface Vacancy alignn_ff 6 44 35 chgnet 0 0 0 eqV2_153M_omat 1 2 0 eqV2_31M_omat 0 0 0 eqV2_31M_omat_mp_salex 0 0 0 eqV2_86M_omat 0 1 0 eqV2_86M_omat_mp_salex 0 0 0 mace 0 1 0 mace-alexandria 0 1 0 mace-mpa 0 0 0 matgl 0 0 0 matgl-direct 0 0 0 mattersim 0 0 0 orb-d3-v2 0 0 0 orb-v2 0 0 0 sevennet 0 0 0
@@
uMLFF Type err_a err_b err_c err_vol err_kv err_c11 err_c44 Time Å Å Å Å3 GPa GPa GPa seconds alignn_ff 0.011 0.011 0.014 0.42 135 196 74 33,081.6 chgnet 0.036 0.038 0.072 3.29 89 59 46 14,102.6 eqV2_153M_omat 0.022 0.025 0.055 2.54 117 75 32 304,173.9 eqV2_31M_omat 0.015 0.016 0.052 2.25 114 78 37 81,420.3 eqV2_31M_omat_mp_salex 0.016 0.017 0.043 1.73 112 46 41 65,971.6 eqV2_86M_omat 0.022 0.025 0.047 2.26 118 86 35 218,168.4 eqV2_86M_omat_mp_salex 0.016 0.017 0.054 2.23 113 43 35 146,318.1 mace 0.027 0.027 0.075 2.85 94 42 38 27,620.4 mace-alexandria 0.068 0.072 0.191 5.65 78 69 45 62,932.5 mace-mpa 0.040 0.042 0.084 3.58 110 35 34 27,255.2 matgl 0.043 0.046 0.106 4.23 83 64 41 10,446.1 matgl-direct 0.043 0.042 0.089 3.00 84 72 50 13,420.6 mattersim 0.027 0.028 0.089 3.11 110 32 32 19,047.3 orb-d3-v2 0.025 0.025 0.032 1.60 130 77 74 15,118.9 orb-v2 0.015 0.016 0.044 1.91 118 81 74 12,184.7 sevennet 0.028 0.030 0.094 3.39 103 42 37 17,739.3
@@
When considering the use of a pretrained uMLFF, the most important considerations should be accuracy, rate of convergence, scalability, and computational cost. After running the full CHIPS-FF workflow for the test set of 104 materials, we performed an analysis of which structural relaxation calculations reached convergence (force-maximum threshold of 0.05 eV/Å within 200 steps) for bulk materials, surfaces and vacancies. This analysis is presented in Table 1 (FrechetCellFilter) and Table S3 (ExpCellFilter), where the percentage of unconverged results is depicted for each calculation type. Aside from ALIGNN-FF, all other uMLFF have a near-perfect convergence rate for bulk structures. With regards to surfaces and defects, the convergence rate is more variable when using the ExpCellFilter, but near-perfect when using the FrechetCellFilter (with the exception of ALIGNN-FF). Interestingly, the OMat24 models significantly vary in terms of convergence for surfaces depending on training data (for the ExpCellFilter). Specifically, the OMat24 models trained on OMat+MPTrj+sAlexandria have a significantly higher rate of convergence when compared to the models solely trained on OMat (see Table S3). These results (Table 1 vs. Table S3) demonstrate the robustness of the FrechetCellFilter in ASE for these uMLFF models.
@@
For each of the 104 structures, we performed structural relaxations, fitting of the energy vs. volume curve (equation of state), calculation of the elastic tensor, calculation of the phonon band structure, relaxation of all nonpolar surfaces (among [1, 0, 0], [1, 1, 1], [1, 1, 0], [0, 1, 1], [0, 0, 1], and [0, 1, 0]) , and relaxation of each type of point defect in the material. The computational timing and errors in the lattice constants, and elastic properties for each uMLFF model are depicted in Table 2 (computed with the FrechetCellFilter). An interactive version of these results is available on the JARVIS-Leaderboard platform (see Fig. 2b) for Pearson Correlation Coefficient metric), where it will be continuously updated as new uMLFF models are developed and released. In addition, structural results computed using the ExpCellFilter are depicted in Table S4. Unsurprisingly, we observe more accurate lattice parameters computed with the FrechetCellFilter due to its robustness in optimizing structural parameters (see Table 2 vs. Table S4). In Table 2, we see that all uMLFF models are able to compute reasonably accurate structural parameters. ALIGNN-FF does an excellent job of simultaneously capturing a, b, and c. This is expected due to the fact that ALIGNN-FF was trained on the JARVIS-DFT dataset (vdW-DF-optB88) and the target/“ground truth”, in addition to the initial structures, are from JARVIS-DFT. The OMat24 models perform exceedingly well, but are the most computationally expensive. The ORB models also perform exceedingly well for structural relaxation and have the advantage of being close to an order of magnitude more computationally efficient. One interesting trend we observe in our results for this set of 104 materials is that the error for lattice constant c is substantially higher than a and b. This can be due in part to the test set containing a significant portion ( 10 ) of vdW bonded (“2D-bulk-like”) layered structures (see Fig. S1). Since most of these uMLFF were trained on PBE data without vdW corrections (with the exception of ALIGNN-FF), we can expect larger errors for vdW systems. For example, we observe significant errors for hexagonal BN (JVASP-62940) when relaxed with orb-v2, where we observe over 10 error in c. These errors are drastically more significant when using the ExpCellFilter along with orb-v2, where we observed a 130 error in c for JVASP-62940 (see Table S4 for total MAE). Fortunately, some of these models such as MACE and ORB have explicit dispersion corrections that can be added to more accurately address vdW interactions 67, 68, 53, 54. To test this, we ran the workflow for dispersion-corrected orb-d3-v2. For orb-d3-v2, we observe a reduction in error for c and volume (see Table 2). For JVASP-62940 relaxed with orb-d3-v2, the percent error in c is reduced to 0.02 . This emphasizes the importance of dispersion corrections in uMLFF architectures for vdW bonded materials.
Fitting the equation of state to obtain bulk modulus and calculating the elastic tensor prove to be a more challenging task for these uMLFF models, emphasizing difficulties in modeling the potential energy landscape far from equilibrium. We found that models such as MACE-Alexandria and MatGL have superior performance in predicting the energy vs. volume curves compared to the OMat and ORB models. For contributions to the elastic tensor (C11 and C44), we also observe relatively high errors. We see that only a few models can simultaneously predict C11 and C44 with reasonable accuracy (MAE less than 50 GPa). Interestingly, we found that eqV2_31M_omat, eqV2_86M_omat and eqV2_153M_omat can predict C44 with reasonable accuracy, but have much higher prediction errors for C11, while eqV2_31M_omat_mp_salex and eqV2_86M_omat_mp_salex do a better job at predicting C11 and C44 simultaneously. We observe that MACE-MPA-0 and MatterSim give the best results for simultaneous C11 and C44 predictions.
| Method | Displacements (Å) | |||
| 0.2 | 0.05 | 0.01 | 0.001 | |
| alignn_ff | 155 | 155 | 157 | 164 |
| chgnet | 70 | 75 | 83 | 91 |
| eqV2_153M_omat | 49 | 47 | 49 | 75 |
| eqV2_31M_omat | 48 | 46 | 53 | 73 |
| eqV2_31M_omat_mp_salex | 48 | 46 | 112 | 219 |
| eqV2_86M_omat | 51 | 47 | 49 | 77 |
| eqV2_86M_omat_mp_salex | 48 | 46 | 103 | 208 |
| mace | 58 | 60 | 60 | 60 |
| mace-alexandria | 82 | 84 | 85 | 85 |
| mace-mpa | 50 | 50 | 50 | 50 |
| matgl | 82 | 93 | 94 | 94 |
| matgl-direct | 86 | 101 | 102 | 101 |
| mattersim | 48 | 47 | 47 | 47 |
| orb-d3-v2 | 53 | 50 | 85 | 181 |
| orb-v2 | 49 | 50 | 108 | 201 |
| sevennet | 54 | 55 | 56 | 56 |
Out of the 104 materials in our test set, there were 84 matching entries for DFT phonon calculations in JARVIS, from which the composite MAE of the phonon band structure was computed in Table 3. We performed these phonon calculations with the uMLFFs at various displacements in the finite-displacement calculation of force constants. We found that models such as ALIGNN-FF and MatGL have relatively large errors in phonon predictions. For certain uMLFFs, we observe a consistent phonon MAE as the displacements get smaller (MACE-MP-0, MACE-MPA-0, MACE-Alexandria, Mattersim, SevenNet). For ORB and OMat models, we see a drastic increase in phonon MAE as the displacements become smaller. This signifies that ORB and OMat models could possibly perform poorly in the low-force regime due to increased noise from the small displacements. This is consistent with recent results, where it was reported that OMat and ORB models had substantial errors in phonon predictions (for small displacements) when compared to models such as MatterSim, MACE, CHGNet, and SevenNet for 10,000 materials 90. This work hypothesized that the errors are due to ORB and OMat computing forces as a direct output of the neural network as opposed to calculating the forces by evaluating the derivative of the energy with respect to the atomic positions (through back-propagation) 90. MatterSim offers similar performance (and consistency across atomic displacements) in terms of predicting the phonon band structure when compared to accurate equivariant uMLFFs such as MACE and SevenNet. This is especially interesting since MatterSim-v1 is not an equivariant uMLFF. In addition, we see improved phonon results for the newer MACE-MPA-0 when compared to MACE-MP-0.
@@
In addition to summing the computational timing for the entire workflow, we also examined the scaling behavior of each uMLFF (on CPU). These results for supercells of Cu (up to 10,000 atoms) are depicted in Fig. 2a). Unsurprisingly, we see that OMat model scales much more drastically with system size as compared to the other uMLFF models. In addition, we found that the invariant models have more favorable scaling than the equivariant models such as MACE and SevenNet, with SevenNet having slightly better scaling than MACE. The scalability of ORB and MatterSim is a huge advantage due to their accuracy of various material property predictions. Similar scalability studies were conducted in Ref. 54 The quantities of surface energy and vacancy formation energy are difficult to predict with machine learning methods and provide a rigorous benchmarking test for uMLFF architectures. Out of the 104 materials, there exists 85 entries for surface DFT calculations and 48 entries for defect DFT calculations within JARVIS-DFT. We used these calculations to compare our uMLFF results. Fig. 3a) and Fig. 3b) depict the parity plots and corresponding MAE for surface energy and vacancy formation energy of our test sets. The very low error for surface energy (0.16 J/m2) and vacancy formation energy (0.36 eV) highlight some of the successes of recent state-of-the-art models such as OMat24, ORB, MACE-MPA-0 and MatterSim. We also observe reasonably accurate results for MACE-MP-0 and SevenNet when predicting vacancy formation energy and surface energies. Recently, the proprietary model PFP from Preferred Networks Inc. was evaluated for our CHIPS-FF surface energy benchmark and achieved an MAE of 0.19 J/m2 124, achieving similar accuracy to the OMat24 and ORB models. The low cost of ORB, high rate of convergence and high accuracy make it a viable tool to relax larger surfaces and defect supercells.
@@
In addition to benchmarking the properties of bulk materials, we utilized the CHIPS-FF workflow for interface calculations. Fig. 3c depicts the parity plot for work of adhesion (Wad) for various material interfaces. The MAE for Wad is computed with respect to “ground truth” data taken from Ref. 123 (from experiment and theory). As seen in Fig. 3c, predictions for Wad are extremely poor for each uMLFF type. The parity plot in Fig. 3c is zoomed in to focus on a subset of calculations, but the MAE is computed for the entire test set of interfaces, which include significant outliers. This is not surprising due to the fact that none of these uMLFF are trained on interface data. We went on to analyze one of the best performing predictions for an Al(111)Al2O3(001) interface (see red box on Fig. 3c). From here, we performed an in-plane (xy) grid search to find the optimal energy configuration of the interface. In addition, this in-plane scan can be a test of how smooth the potential energy surface is. On the right panel of Fig. 3c we see the results of this interface scan for a number of uMLFFs. We observe a smoother potential energy surface for MACE, SevenNet, MatterSim and OMat24 models. Although the MAE is extremely high for all uMLFF models for Wad, it is the lowest for ALIGNN-FF. This can be due in part to ALIGNN-FF being trained on vdW-corrected DFT calculations and having a substantial amount of exfoliable materials in the dataset.
@@
In addition to testing these uMLFF models on a smaller and more focused dataset, we also tested on a larger and more diverse dataset for force predictions. Unlike total energy, forces are more independent of exchange-correlation functional and the accuracy of force predictions on various datasets can give valuable insight into how a particular model performs. We chose to benchmark each uMLFF on the MLEARN dataset from Ref. 125, which consists of face-centered cubic (Cu, Ni) and body-centered cubic (Li, Mo) metals and diamond group IV semiconductors (Si, Ge) which span a wide range of crystal structures and bonding environments ( 200 - 300 data points for each element). These results are depicted in Table 4. In addition, we benchmarked the accuracy of force predictions on very large datasets that were used to train uMLFFs. These datasets included ALIGNN_FF_DB (307,000 used to train ALIGNN-FF), MPF (188,000 used to train M3GNet), and MPTrj (1.58 million used to train CHGNet, MACE, SevenNet and used in the training of ORB and OMat models). These results are depicted in Table 5. Unsurprisingly, superior accuracy is obtained across all datasets when the OMat and ORB models are used to predict forces. We also observe highly accurate forces for MatterSim and a substantial reduction in force error when going from MACE-MP-0 to MACE-MPA-0. The CHIPS-FF package has built in functions to perform these benchmarking calculations on each respective dataset, with a flexible framework to add additional datasets (such as OMat and Alexandria) and new uMLFFs as they develop over time.
@@
ALIGNN_FF_DB (307k) MPF (188k) MPTrj (1.58M) MLFF Type Error (eV/Å) Time (s) Error (eV/Å) Time (s) Error (eV/Å) Time (s) alignn_ff 0.472 2.9 x 104 0.639 4.1 x 104 0.581 3.5 x 105 chgnet 0.088 3.2 x 104 0.101 1.1 x 105 0.060 7.6 x 105 eqV2_153M_omat 0.050 1.5 x 106 - - - - eqV2_31M_omat 0.051 3.0 x 105 0.046 3.5 x 105 - - eqV2_31M_omat_mp_salex 0.055 3.9 x 105 0.032 4.3 x 105 - - eqV2_86M_omat 0.052 9.2 x 105 0.043 9.3 x 105 - - eqV2_86M_omat_mp_salex 0.056 9.2 x 105 - - - - mace 0.104 5.0 x 104 0.125 1.8 x 105 0.059 1.1 x 106 mace-alexandria 0.221 1.3 x 105 0.395 2.7 x 105 - - mace-mpa 0.072 7.0 x 105 0.061 1.3 x 105 0.039 1.9 x 106 matgl 0.091 1.1 x 104 0.066 3.4 x 104 0.100 1.4 x 105 matgl-direct 0.104 1.7 x 104 0.095 4.2 x 104 0.118 2.7 x 105 mattersim 0.071 8.7 x 104 0.047 7.1 x 104 0.059 1.7 x 106 orb-v2 0.062 5.0 x 104 0.049 8.4 x 104 0.028 7.5 x 105 sevennet 0.092 4.4 x 104 0.081 6.0 x 104 0.041 5.0 x 105